# Chameleon Channels: Measuring YouTube Accounts Repurposed for Deception and Profit

### v1

@@ Abstract
Online content creators spend significant time and effort building their user base through a long, often arduous process, which requires finding the right “niche” to cater to. So, what incentive is there for an established content creator known for cat memes to completely reinvent their page channel and start promoting cryptocurrency services or cover electoral news events? And, if they do, do their existing subscribers not notice?
We explore this problem of repurposed channels, whereby a channel changes its identity and contents. We first characterize a market for “second-hand” social media accounts, which recorded sales exceeding USD 1M during our 6-month observation period. By observing YouTube channels (re)sold over these 6 months, we find that a substantial number (37%) are used to disseminate potentially harmful content, often without facing any penalty. Even more surprisingly, these channels seem to gain rather than lose subscribers.
To estimate the prevalence of channel repurposing “in the wild,” we also collect two snapshots of 1.4M quasi-randomly sampled YouTube accounts. In a 3-month period, we estimate that $\sim$0.25channels—collectively holding $\sim$44M subscribers—were repurposed. Through a set of experiments, we confirm that these repurposed channels share several characteristics with sold channels—mainly, the fact that they had a significantly high presence of potentially problematic content. Across repurposed channels, we find channels that became disinformation channels, as well as channels that link to web pages with financial scams. We reason that abusing the residual user trust placed on these channels is advantageous to financially- and ideologically-motivated adversaries. This phenomenon is not exclusive to YouTube and we posit that the market for cultivating organic audiences is set to grow, particularly if it remains unchallenged by mitigations, technical or otherwise.
@@ 1 Introduction
Online content creators and influencers are central to digital spaces. Teens and adults in the United States alike consume hours of content every day from their favorite creators [1, 2]: from news [3], and recommendations for products and services [4] to advice on financial, professional, and medical topics [5]. Being a content creator is a highly sought after profession, with more than 57% GenZers saying they would like to become an influencer if given the opportunity [6]. In that context, it is unsurprising that tools and services dedicated to growing one’s online influence abound. Offerings range from cheap bulk bot-based engagement to increasingly more sophisticated products involving engagement from real humans.
A new product type has gained popularity in recent years: the sale of credentials to an account with a pre-established audience and reputation. A recent study found over 38,000 accounts and an advertised value of over USD 64 million [7]. According to Beluri et al. [7], Fameswap boasts the highest number of sellers and has experienced an almost 9x growth since being first described by Chu et al. in 2022 [8].
@@
A key question is whether platforms like Fameswap constitute a new paradigm for online engagement—as opposed to just a new packaging for already known artificial engagement strategies. If these ready-made accounts represent a new type of product, what are the implications for the platforms whose accounts are commercialized? We aim to answer these questions by observing how these accounts are used after being sold. We conduct this observation on YouTube by leveraging the persistent nature of channel IDs—invariant identifiers that map to a channel, and do not change even if the channel entirely alters its identity by changing handles, title, description, and/or content–––what we call repurposing a channel, shown in Figure 1.
@@
We hypothesize that repurposing channels alleviates a critical constraint for financially and ideologically motivated adversaries: rapid access to distribution without the resource expenditure required to cultivate an audience from scratch. To evaluate this hypothesis, we explore the following research questions:
  • RQ1: How are channels sold in online marketplaces repurposed?
  • RQ2: Do repurposed channels disseminate potentially problematic content?
  • RQ3: To which extent are channels repurposed on YouTube?
  • RQ4: What are indicators of channel reuse?
To investigate these questions, we conducted a mixed-methods study in which we characterize the Fameswap marketplace and collect YouTube channels advertised for sale, exploring the time it takes until a channel is repurposed, the change in subscribers after the change, and the rate of repurposing. We then collect a large quasi-random sample of YouTube channels from an analytics platform named Social Blade, and identify repurposed channels in the wild. Using our Fameswap and Social Blade samples of repurposed channels, we annotated each channel based on potentially problematic content categories.
Our work provides an in-depth view of an emerging vector that currently faces few mitigations. Its implications are the same for all other platforms for which there are accounts for sale.
@@ 2 Related Work
Abusing Reputation Systems. Online reputation refers to metrics that signal user trustworthiness. First developed for online marketplaces, reputation systems—reviews, ratings, etc.—were key to their success [9]. These systems were later adopted by social media to incentivize contributions [10], encourage participation [11, 12], and signal authenticity [13]. Today, they take the form of likes, followers, badges, and similar indicators.
These indicators are valuable targets for manipulation. Over the past two decades, researchers have documented widespread abuse across platforms [14, 15, 16, 17, 18]. Misuse has evolved from bots and compromised accounts [19, 20, 21, 22] to more human-like tactics: crowdsourced engagement [23, 24, 25], automation of real user accounts [26], collusion rings [27, 28], and click farms [29, 30].
Account hijacking has similarly shifted toward high-value targets. Rather than compromising many small accounts, actors now focus on fewer but influential ones [31]. This shift is tied to the ease of monetizing influence, particularly in finance and crypto spaces. Notable examples include the SEC’s hijacked account used for market-moving tweets [32] and Vitalik Buterin’s account used to spread phishing links [33].
@@
Licit Yet Harmful Online Influencers. Influential accounts that promote potentially harmful content but within gray areas have become an increasingly prevalent issue. For example, many reported cases of influencers promoting “meme coins” have left viewers with little recourse after the coin collapses [35, 36]. It is unclear whether YouTube’s terms-of-service restrict this behavior, whose (il)legality is not even completely settled. Similarly, influential accounts have also been used to disseminate problematic narratives, such as smear campaigns, before elections. The latest case occurred just last year, when a criminal investigation linked payments from a Russian disinformation campaign to American influencers [37]. This example also illustrates the increasing interplay between financially motivated actors (in this case, influencers) and ideologically motivated actors (state sponsors).
Disinformation and Underground Markets. Services to assist or implement disinformation campaigns have often overlapped with underground markets and communities [38, 39, 40, 41]. On the one hand, these marketplaces specialize in the dissemination of content that is frequently at odds with platform guidelines, for example, advertising restricted products (e.g., pharmaceuticals, tobacco products, unregulated gambling websites, etc.), including disinformation. On the other hand, these marketplaces also provide tools and infrastructure that can aid disinformation operators (e.g., coordinated engagement, bulletproof hosting, DDoS protection, advertising networks, etc.) [42]. Most importantly, off-platform sources of monetization are an important driver for channels with problematic content to thrive [8, 43, 44].
Domain and Handle Reuse. Our work shares many similarities with work on residual trust across domains and namesquatting [45, 46]. Most related to our work, Mariconti et al. study the reuse of profile names on on Twitter [47]. Their study focuses on accounts that register handles that had been de-registered to capture the residual reputation and backlinks pointing to those handles [47]. However, while some of our findings, such as prevalence of issues and the fact that reuse is often used for questionable purposes [47], overlap, our work is substantially different, given that a repurposed channel (YouTube) carries all previous followers. The market for ready-made accounts has remained largely understudied, only receiving a brief overview by Chu et al. [8], and more recently a more in-depth characterization by Beluri et al. [7]. To our knowledge, our work is the first to explore how sold accounts are repurposed and to identify channel repurposing in the wild.
@@
Figure 1: On the left, a channel listed for sale on Fameswap about entertaining facts with 1.19M subscribers. Videos are created predominantly with AI tools. This channel would later change to a news channel discussing political issues with no trace of its previous identity (on the right). For privacy, the channels’ handles and titles are blurred.
@@ 3 Background
We provide background on the market we study and relate it to some of the earlier findings in the literature.
3.1 The Fameswap Market

### v2

@@ Abstract
Online content creators spend significant time and effort building their user base through a long, often arduous process that requires finding the right “niche” to cater to. So, what incentive is there for an established content creator known for cat memes to completely reinvent their page channel and start promoting cryptocurrency services or covering electoral news events? And, if they do, do their existing subscribers not notice?
We explore this problem of repurposed channels, whereby a channel changes its identity and contents. We first characterize a market for “second-hand” social media accounts, which recorded sales exceeding USD 1M during our 6-month observation period. Observing YouTube channels (re)sold over these 6 months, we find that a substantial number (53%) are used to disseminate policy-sensitive content, often without facing any penalty. Even more surprisingly, these channels seem to gain rather than lose subscribers.
We estimate the prevalence of channel repurposing “in the wild,” using two snapshots of $\sim$1.4M YouTube accounts sampled from an ecologically valid proxy. In a 3-month period, we estimate that $\sim$0.25% channels were repurposed. Through a set of experiments, we confirm that these repurposed channels share several characteristics with sold channels—mainly, the fact that they have a significantly high presence of policy-sensitive content. Across repurposed channels, we find channels similar to those used in influence operations, as well as channels used for financial scams. Repurposed channels have large audiences; across two observed samples, repurposed channels collectively held $\sim$193M and $\sim$44M subscribers. We reason that purchasing an existing audience and the credibility associated with an established account is advantageous to financially- and ideologically-motivated adversaries. This phenomenon is not exclusive to YouTube and we posit that the market for cultivating organic audiences is set to grow, particularly if it remains unchallenged by mitigations, technical or otherwise.
@@ 1 Introduction
In 2022, more than 22,000 users suddenly found themselves following @HonorColoradoOk—an account affiliated with Paraguay’s ruling political party—despite never opting in [4]. Overnight, timelines were flooded with political content, just a year ahead of the presidential elections. The account had previously operated as @InfoVacunatePy, a government-linked source for vaccination site updates, but was allegedly sold and repurposed. The abrupt change in identity, amplified by the polarized political landscape, quickly drew user complaints and media attention. Within days, the owners removed the account, but not before sparking public outcry and exposing how easily an established audience can be redirected toward new—sometimes questionable—content. Yet, in many cases, these identity changes go unnoticed.
The ability to repurpose accounts has long enabled the buying and selling of social media identities, historically conducted through private exchanges on forums and chat applications. These transactions were opaque and difficult to systematically study. In recent years, however, the rise of dedicated marketplaces combined with the reduced costs of content creation afforded by AI has dramatically expanded both the scale and visibility of this ecosystem. A recent study documented 38,000 accounts for sale, with an advertised value exceeding USD 64 million [12]. Fameswap, one of the most active marketplaces, has now ninefold more active listings since its initial observation by Chu et al. in 2022 [15].
@@
A key question is whether platforms like Fameswap constitute a new paradigm for online engagement—as opposed to just a new packaging for already known artificial engagement strategies [60]. If these ready-made accounts represent a new type of product, what are the implications for the platforms whose accounts are commercialized? We address this question by observing how these accounts are used after being sold. We examine this question on YouTube by leveraging two measurement advantages. First, the rise of increasingly transparent account marketplaces provides direct visibility into accounts offered for sale. Second, YouTube assigns each channel a persistent identifier that remains fixed even when the channel changes its handle, title, description, or content, allowing us to track what we call repurposing, illustrated in Figure 1.
@@
Repurposing channels may alleviate a critical constraint for financially and ideologically motivated adversaries: rapid access to content distribution without the resource expenditure required to cultivate an audience from scratch. We explore this possibility through three exploratory questions:
  • RQ1: How are channels used after being repurposed on YouTube?
  • RQ2: How prevalent is repurposing on YouTube?
  • RQ3: What are indicators of repurposed channels?
We then test the following hypothesis:
  • H1: Repurposed accounts (either advertised for sale or obtained in the Social Blade sample) disseminate policy-sensitive content at significantly higher rates than non-repurposed accounts.
To investigate these questions, we collect YouTube accounts advertised for sale from Fameswap, one of the largest social media account marketplaces [12], and from Social Blade, an analytics platform that tracks 68M+ YouTube accounts (§ 4.1). We then define and annotate channel repurposing and presence of policy-sensitive content (i.e., content which, due to its sensitive nature and potential for abuse, is regulated by YouTube’s policies [100]) in §§ 4.24.3. In § 5, we explore the impact of repurposing on growth, suspension, and content changes through repurposed Fameswap accounts. Then, by labeling repurposing on the Social Blade sample, we estimate the prevalence of repurposing across YouTube in § 6.1. Lastly, we compare repurposing indicators and differences in content by comparing repurposed channels (both from Fameswap and Social Blade) against a non-repurposed control group (§ 6.2).
We find that repurposed accounts continue to accrue subscribers after repurposing, that most of them are not suspended, and that repurposing is prevalent. We estimate that 0.25% of all accounts indexed by Social Blade were repurposed between January and March 2025 and argue that this estimate generalizes to YouTube as a whole. We find that repurposed channels post higher rates of policy-sensitive content, especially ideological and financial content. During our study, we observed that repurposed Fameswap accounts had 193M+ subscribers and repurposed Social Blade accounts had 43M+ subscribers. Purchasing and repurposing accounts seems to be a viable way to quickly distribute content which may affect the integrity of online platforms, particularly if the content may result in financial harm to users or is part of a political influence operation. We argue that these findings generalize to other platforms where accounts are traded on secondary markets and where modifying profile indicators (e.g., handles or usernames) is trivial and difficult to detect.
@@ 2 Related Work
Abusing Reputation Systems. Online reputation refers to metrics that signal trustworthiness. Reputation systems—reviews, ratings, etc.—were first developed in online marketplaces and instrumental to their success [75]. These systems were later adapted by social media to incentivize various behaviors [67, 48, 9] and signal authenticity [80]. Today, they take the form of likes, followers, badges, and similar indicators. These indicators are valuable targets for manipulation. Researchers have documented widespread abuse across platforms [57, 54, 94, 73, 60]. Misuse has evolved from bots and compromised accounts [71, 78, 23, 14] to more human-like tactics: crowdsourced engagement [86, 95, 59], automation of real user accounts [18], collusion rings [87, 27], and click farms [33, 20].
Account hijacking has similarly shifted toward high-value targets. Rather than compromising many small accounts, actors now focus on fewer but influential ones [24]. This shift may be tied to the ease of monetizing influence, particularly in finance and crypto spaces. Notable examples include the SEC’s hijacked account used for market-moving tweets [79] and Vitalik Buterin’s account used to spread phishing links [47].
@@
Licit Yet Harmful Online Influencers. Influential accounts that promote potentially harmful content but within gray areas have become an increasingly prevalent issue. For example, many reported cases of influencers promoting “meme coins” have left viewers with little recourse after the coin collapses and they incur losses [16, 46]. It is unclear whether YouTube’s terms of service prohibit influencers from promoting speculative financial instruments such as meme coins, in part because the legality of such promotions remains unsettled. Nonetheless, their financial impact on viewers is clear. Similarly, influential accounts have also been used to disseminate problematic narratives, such as smear campaigns, during election periods. The latest case occurred just last year, when a criminal investigation linked payments from a Russian disinformation campaign to American influencers [32]. This example also illustrates the increasing interplay between financially motivated actors (in this case, influencers) and ideologically motivated actors (state sponsors).
@@
Figure 1: On the left, a channel listed for sale on Fameswap about entertaining facts with 1.19M subscribers. Videos are created predominantly with AI tools. This channel would later change to a news channel discussing political issues with no trace of its previous identity (on the right). Note, the channel on the left has been reconstructed using metadata. To not drive traffic to the channels, handles and titles are blurred.
@@
Disinformation and Underground Markets. Services to assist or implement disinformation campaigns have often overlapped with underground markets and communities [85, 34, 55, 66]. On one hand, these marketplaces specialize in the dissemination of content that is frequently at odds with platform guidelines, for example, advertising restricted products (e.g., pharmaceuticals, tobacco products, unregulated gambling websites, etc.), including disinformation. On the other hand, these marketplaces also provide tools and infrastructure that can aid disinformation operators (e.g., coordinated engagement, bulletproof hosting, DDoS protection, advertising networks, etc.) [36]. Most importantly, off-platform sources of monetization are an important driver for channels with problematic content to thrive [15, 40, 10].
Domain and Handle Reuse. Our work shares many similarities with work on residual trust across domains and namesquatting [69, 50]. Most related to our work, Elmas et al. study account misleading repurposing on Twitter, whereby an account changes profile attributes to use the account for a new purpose while retaining its followers [25]. They propose a supervised learning method for detection and find common behavioral patterns across repurposed accounts (e.g., repurposing after inactivity or tweet deletion [25]. Our work builds upon theirs but extends it by estimating the extent to which repurposing occurs across a whole platform and comprehensively quantifying the types of content that are posted after repurposing. Also related is work by Mariconti et al. who study the reuse of profile names on on Twitter [56]. Their study focuses on accounts that register handles that had been de-registered to capture the residual reputation and backlinks pointing to those handles [56]. In our case, repurposed channels carry all previous followers.
@@ 3 Background
Fameswap is an online marketplace for buying and selling accounts on YouTube, TikTok, Instagram, Twitter/X, and websites. Recent measurements identify it as the largest such market by number of sellers [12]. The site resembles a standard e-commerce platform: each account has a dedicated listing with a description, price, seller reviews, and a content category (e.g., humor, sports). Buyers can bid, and the platform offers escrow, dispute resolution, and account verification—features similar to marketplaces like eBay. Fameswap earns revenue from escrow fees (3% or a USD 50 minimum) and premium accounts, which offer advanced search, additional metrics, lower fees, and more. Payments are accepted via PayPal, wire transfers, and cryptocurrencies (e.g., BTC, ETH, USDC). Fameswap verifies account ownership by providing sellers with a unique randomized string that should be placed in the advertised social media profile. For example, a seller receives the string aSgc5H3s and places it temporarily in their YouTube channel description, allowing Fameswap to verify the ownership, similar to using TXT records to verify domains.
Fameswap listings link the corresponding channel’s YouTube ID. Different from channel titles, handles (“@SomeChannel”) are unique and can be accessed through “www.youtube.com/@SomeChannel [77].” Handles resolve to a unique channel ID, allowing a channel to change its handle but keep pointing to the same channel. YouTube’s channel ID is a string that begins with “UC”, followed by 22 characters (letters, numbers, dashes, and underscores). Channel IDs can also be used to access a channel in the form: “www.youtube.com/channel/UC…”. While an account can change its handle, its channel ID cannot change, which allows us to monitor accounts over time.
The market for ready-made accounts (as opposed to the market for artificial engagement [60]) has remained largely understudied, only receiving a brief overview by Chu et al. [15] in their study of illicit alternative monetization strategies on YouTube. In their study, they identified five marketplaces for social media accounts: SWAPD[74], Accs-Market.com [5], Trustiu (now inactive), ViralAccounts [83], and Fameswap [26]. Beluri et al., expand on Chu et al.’s work by conducting a more comprehensive financial characterization of social media account markets [^1] Their study became public after our data collection began and is referenced retrospectively.[12]. Their results indicate that Fameswap is a leading marketplace and thus an appropriate candidate for study [12]. In addition to the platforms described in these studies, we note that a substantial amount of trade occurs in other platforms that are not described, such as the forum OGUser (similar to SWAPD) and Telegram channels.
4 Methods
We measure the repurposing of YouTube accounts and their participation in disseminating policy-sensitive content through the lens of sold accounts extracted from a social media marketplace (Fameswap) and by observing handle changes and channel repurposing in the wild from an ecologically valid proxy (Social Blade). We structure our study in three parts, as depicted in Figure 3. First, we define and characterize channel repurposing across Fameswap channels (§ 4.2). Second, we develop a codebook for topic extraction and use it to annotate repurposed Fameswap channels and observe content dissemination after repurposing (§ 4.3). Using these annotations in § 5, we characterize the Fameswap marketplace and the repurposed accounts obtained from there. Third, we observe repurposing in the wild and estimate its prevalence. We then use a framework for deriving statistically valid estimates using textual annotations and quantitatively test what features are indicative of repurposing across samples (§ 6).
4.1 Data Collection and Sources
To carry out all our experiments, we collected, labeled, and categorized YouTube metadata from two sources, Fameswap and Social Blade, as shown in Figure 2. We provide a summarized description of our samples, as well as instructions for data access and replication in Appendix B. An additional investigation using the Wayback Machine is discussed in Appendix C.1. Lastly, ethical considerations are discussed in Appendix A.
4.1.1 YouTube Channels For Sale on Fameswap
As illustrated in Figure 2(a), we conducted daily scrapes of Fameswap from October 26, 2024, to March 31, 2025. The Fameswap interface provides a paginated list of all account listings. We scraped all historical listings. This set includes all listings not deleted or hidden from the website. Using the YouTube IDs linked to each listing, we conducted regular scrapes of all collected channels. We scraped channels every 3 days on average. As of March 31, 2025, we collected 4,641 YouTube channel IDs advertised for sale, each with about eight observations per month. Form Fameswap listing, we scraped their YouTube channel using the YouTube Data API v3.
4.1.2 Collecting a Large Ecologically Valid Sample of YouTube through Social Blade
To estimate the prevalence of channel repurposing and identify repurposing indicators beyond marketplaces, we aimed to observe this phenomenon “in the wild.” Randomly sampling YouTube is difficult due to the lack of a centralized channel directory and method to enumerate channel IDs [39]. Horta Ribeiro and West address this by sampling from Social Blade [38], an analytics site that has indexed over 68M channels in the past 17 years [70]. Social Blade is a widely used reference for creators [89] and has informed multiple YouTube studies [65, 28]. While Social Blade’s crawling process is not public, it likely favors more popular channels. This, however, is acceptable given that our focus is content creators with larger audiences.
To conduct this study, we obtained a dataset of YouTube IDs collected by Horta Ribeiro and West. We randomly sampled 1.4M channels and conducted an initial scrape from December 23, 2024, to January 21, 2025 (Figure 2(b)). We collected metadata from 1,397,586 channels and 139M corresponding videos. A second scrape, from March 21–31, 2025, captured updated metadata and 20M new videos. Only 1,351,912 channels returned data (a 3.3% deletion rate). We used yt-dlp, a common tool for archiving YouTube content [91, 62, 64]. While yt-dlp is commonly used in academic projects, we discuss potential ethical implications in Appendix A.
4.2 Labeling Repurposed Channels
Our first goal is to detect channels that have been repurposed, such as the example shown in Figure 1. In this figure, we observe how a channel that was posting primarily entertainment facts then becomes a channel that shares political news. Repurposing a channel allows an individual to portray herself as legitimate, distribute content to an existing audience, or potentially increase the chances that the recommendation algorithm amplifies new content [96, 52]. We first apply a qualitative approach to define what constitutes a channel being repurposed. We use these qualitative insights to develop an LLM prompt which we can then use to scale our annotations.
Determining whether a channel has been repurposed is an entity matching problem. We can say that a tuple representing a paired observation in time of a given channel constitutes a channel being repurposed if both observations do not refer to the same identity. However, given that we do not have ground truth for channel identities, we first create a definition based on human perception and then enhance this procedure using an LLM. Because LLMs have been trained on large text corpora, LLMs are well-suited for entity matching problems because they can identify connections between entities a human annotator may miss [61].
To create a definition, we followed a three-stage process. First, the lead author manually monitored a random sample of 104 channels (10%) that changed handles after being listed for sale. Each channel was visited weekly over a month to observe changes in content, metadata, and activity. We created standardized weekly snapshots capturing each channel’s videos, thumbnails, titles, timestamps, and descriptions (Appendix B.2). Second, using open coding, two researchers independently analyzed a new random sample of 54 channels (5%) using these documents. Each annotator coded the channels following the prompt: “was the channel repurposed?” Coders discussed their annotations to identify edge cases and refine definitions. Disagreements arose around dormant channels, subtle content shifts, and stylistic similarities with different themes (e.g., channels that had same style of content but around different themes). This coding was exploratory, thus we did not compute agreement scores. Coders prioritized changes in handles and titles, using descriptions and video content as secondary evidence. Minor changes (e.g., @HealthReporter to @TheHealthReporter) were not considered repurposed. Significant discrepancies (e.g., @HealthReporter to @247News) prompted checks for overlap in topics, URLs, or referenced entities. If no continuity was found, the channel was flagged as repurposed. A substantial shift in video content (e.g., religious to crypto content in another language) served as a confirmatory signal but was not sufficient on its own. In the third step, the lead author and an external coder with a large social media following and understanding of the content creation ecosystem reviewed another 54 channels (5%) as a sanity check.
4.3 Labeling Policy-Sensitive Content
Our second goal is to characterize the uses and understand motivations by observing the content disseminated after repurposing a channel. We define as policy-sensitive, content that due to their sensitive nature (e.g., financial, health, political content, etc.) and potential for abuse and harm, is regulated by YouTube’s community guidelines and moderation practices [100, 31]. Informed by prior work [^2] Scholars studying harmful content strongly advise researchers “not to accept how platforms frame and define issues” [82]. This is because content guidelines are often broad and cover a wide range of topics, and definitions often lack specificity [82]. and our initial rounds of annotation, we focus on two main broad types of policy-sensitive content: financial and ideological. Additionally, also due to their prevalence in initial annotations, we labeled channels that contained AI-generated content, copyright infringement, content oriented towards children, and the presence of alternative forms of monetization (all of which are also governed by YouTube’s policies, and hence policy-sensitive).
4.3.1 Annotation for Codebook Development
To develop our category annotation codebook, we adopted an empirical approach. We selected a set of restricted themes from YouTube’s Community Guidelines and refined them in various coding rounds into narrower operational definitions. We discuss each of the content categories and then describe how we used an LLM to scale our annotations. Lastly, we explore the categories of content posted before and after repurposing.
We conducted three rounds of content annotation. First, the lead author selected 10% of all repurposed channels (nn=102) and created an initial codebook identifying which YouTube policies applied to the videos (based on metadata) contained in each channel. Using this initial codebook, the lead author and second author independently coded 54 repurposed Fameswap channels (5% of reused channels). Both authors have extensive experience with the YouTube ecosystem. Each researcher received a summary text document containing all of the channel’s observation snapshots. We provide an example of this document in Appendix B.2. After the first round of coding, the coders discussed the procedure and tweaked the codebook to address the shortcomings. We also extracted additional codes from the open-ended field if the content was governed by community guidelines and not yet captured by existing codes.
The lead author and a member external to the research team with extensive knowledge about the creation of social media content did another round of coding. They coded a new randomly sampled but non-overlapping 5% of the dataset. The goal was to identify and reduce ambiguity in the existing definitions. Together with this external member, we derived our final codebook. The final codebook is found in Appendix C.2. Note that we did not compute agreement during these steps, since the goal of these steps was to iteratively create a codebook, not yet to use the codes for downstream analyses. We address the statisical soundness of our annotations in § 4.4. A discussion of ethics, as related to the annotation procedure, is found in Appendix A.
4.3.2 Content Categories
Ideological Content. Based on YouTube’s misinformation guidelines [104, 103, 101], we created three codes covering political, medical, and news content. News and political content often overlap (i.e., politcal news), but not always. We do not attempt to verify whether videos’ content constitute misinformation or disinformation—the latter defined by its intentionality. We do not verify the veracity of claims, as we discuss in § 7.1. However, the dissemination of political or news-related information through social media channels not associated to established news outlets is a hallmark of influence operations, regardless of the claims’ veracity [92].
In addition to political and news content, we noticed a substantial number of religious videos in our sample. Although religion is not a topic captured in community guidelines, we opted to include it as a code given its frequent co-occurrence with geopolitical content and frequently observed content involving religious leaders with political appointments. Past work has also found religious videos to frequently co-occur with extremist and hateful content [6, 7]. Lastly, we included codes for extremist content and manosphere content, based on past work on radicalization pathways on YouTube [38].
Financial Content. On the financial side, we created a code for Gambling content, which is directly regulated by the guidelines [98]. However, we note a substantial amount of content related to making money online through a variety of ways: pay-per-click websites, dropshipping, trading stocks, cryptocurrencies, get-rich-quick schemes, and ironically, courses to grow and monetize social media accounts. These topics are only partially covered by YouTube’s guidelines (e.g., get-rich-quick schemes, fraud, broadly “spam”, and broadly “scams”) [99]. Cryptocurrency-related content was particularly prevalent. Informed by research on cryptocurrency scams and their potential to harm users financially [45, 53], we create a cryptocurrency code. We capture the rest of the aforementioned content under money-making content.
Other Policy-Sensitive Content. We also identified other types of content in our annotation that are covered by YouTube’s guidelines, such as content that may infringe copyright [106] and AI-generated content [102]. AI-generated content is of particular importance given the raise of “faceless YouTube channels” and their adoption by scammers and spammers [19]. Lastly, we also identified content that appears to be geared towards children (e.g., content involved early childhood cartoons, nursery rhymes, and didactic activities like coloring) and pre-teens (e.g., content involved Roblox, Minecraft, and young influencers). We also labeled this content given YouTube’s policies for content targeting minors and child safety [97]. Lastly, we wanted to explore whether repurposed channels had a higher rate of alternative monetization strategies. Past work has found that alternative monetization strategies help channels dedicated to problematic content stay afloat  [40]. Furthermore, we hypothesized that some forms of alternative monetization strategies, commonly discussed in underground forums [15], may be more associated with repurposing practices.
4.4 Scaling Annotations with LLMs
We use LLMs to scale our repurposing and content annotations for two reasons: (1) estimating repurposing across all Social Blade channels and (2) testing for content differences across samples. In the former case, we can use false positives/negatives to directly correct our estimate. For the latter, we employ a regression model. However, directly using LLM-generated labels in downstream regressions can introduce measurement error—mismatch between predicted and true labels [22]. Information on how to access the prompts is discussed in Appendix B.2.
To correct for this measurement error, Egami et al. introduce a framework for statistically valid inference with LLMs: Design-based Supervised Learning (DSL) [22]. The process involves: (1) predicting labels with an LLM, (2) sampling a subset for expert annotation, and (3) combining predictions and gold-standard labels in DSL regression. This method requires that expert-coded samples have known, non-zero sampling probabilities. This approach allows us to use mall-scale qualitative annotations with large-scale LLM predictions, using the latter as “surrogate” classifiers.
We annotated our data using gpt-4o-2024-11-20 (state-of-the-art at the time of writing) with temperature zero and top $p$ of one. We created LLM prompts based on our final codebooks and definitions following best practices suggested by Ziems et al. [107]. Although there is still a lack of consensus on the optimal parameter choice for LLMs, recent work recognizes these parameters as the current standard [63]. We validate our labels by comparing against expert annotations (also called “gold-standard labels”) and describe the framework we employ to obtain statistically valid estimates using these annotations.
Repurposing Labels. Using a prompt based on our repurposing definition, we annotated the Fameswap and Social Blade channels that we collected. We create a gold-standard dataset composed of a random 10% subsample of human annotated channels for each sample. In our Social Blade sample, this annotation allows us to directly estimate the total set of repurposed channels between January and March 2025, using the false positive/negatives rates to derive a 95% confidence interval. We present the results of this estimation in § 6.1.